I am having trouble finding the correct formula for solving this word problem:
Marie is 3 years older than Nicole and 5 years younger than Anna. In 5 years, twice Anna's age taken from 4 times Nicole's age is 2 years younger than Marie's age. How old are they now?

Answer by lwsshak3(11628)


Marie is 3 years older than Nicole and 5 years younger than Anna. In 5 years, twice Anna's age taken from 4 times Nicole's age is 2 years younger than Marie's age. How old are they now?
***
let x=nicole's age
x+3=marie's age
x+3+5=x+8anna's age
..
In 5 yrs:
x+5=nicole's age
x+3+5=x+8=marie's age
x+8+5=x+13=anna's age
..
4(x+5)-2(x+13)=x+8-2
4x+20-2x-26=x+6
x=12
x+3=15
x+8=20
..
nicole's age=12
marie's age=15
anna's age=20